# Column reduced digital nets

### v1

@@ 1.1 The problem setting
where ${\bm{x}}_{k}=(x_{k}^{(1)},\ldots,x_{k}^{(s)})^{\top}$ are column vectors corresponding to the points used in the QMC rule. Problems of this kind particularly arise in some important applications in statistics and uncertainty quantification. For instance, this approach can be used when approximating the expected value of a function with a multivariate normal random variable with some given covariance matrix, or when approximating the expected value of the solution of a PDE with random coefficients, see, e.g., [4].
@@ 4.2 Error analysis
In many applications of quasi-Monte Carlo, one considers so-called weighted function spaces such as weighted Sobolev or weighted Korobov spaces (see, e.g., [3, 5, 6]). The idea of studying weighted function spaces goes back to the seminal paper [11] of Sloan and Woźniakowski. The motivation for weighted spaces is that in many applications, different coordinates or different groups of coordinates may have different influence on a multivariate problem. To give a simple example, consider numerical integration of a function $f:[0,1]^{s}\to{\mathbb{R}}$, where
@@ 6 Conclusion
Data availability statement
The numerical tests in Section 5 were performed with source code that is publicly available at https://github.com/Vishnupriya-Anupindi/ReducedDigitalNets.jl.
Ethics approval
Since this is a paper on theoretical mathematics, no ethical issues arose during the writing of this paper.

### v2

@@ 1.1 The problem setting
[…]
@@ 2.1 Column reduction for $(t,m,s)$-nets
2.2 Column reduction for $(t,m,{\bf{\textit{e}}},s)$-nets
In [13] Tezuka introduced the concept of $(t,m,\bm{e},s)$-nets, which are a generalization of $(t,m,s)$-nets. In this section, we briefly look at the quality parameter of column reduced nets under this generalized definition of nets. However, for the rest of the paper, we shall then stick to the notion of $(t,m,s)$-nets again.
Definition 4.
Let $\bm{e}=(e_{1},\dots,e_{s})$ and $\bm{d}=(d_{1},\dots,d_{s})$ be integer vectors with $e_{i}\geq 1$ and $d_{i}\geq 0$ for $i\in\{1,\ldots,s\}$, where $s\geq 1$ is the dimension. Let $t,m$ be non-negative integers with $0\leq t\leq m$. A point set ${\mathcal{P}}\subset[0,1)^{s}$ with $b^{m}$ points is called a $(t,m,\bm{e},s)$-net in base $b$ if every elementary interval $J\subseteq[0,1)^{s}$ of volume $b^{t-m}$ and of the form
$J=\prod_{j=1}^{s}\Big{[}\frac{a_{j}}{b^{e_{j}d_{j}}},\frac{a_{j}+1}{b^{e_{j}d_{j}}}\Big{)}$
contains exactly $b^{t}$ points of ${\mathcal{P}}$, where $0\leq a_{j}<b^{e_{j}d_{j}}$ for $j\in\{1,\ldots,s\}$ and $\bm{d}$ satisfies the equation $e_{1}d_{1}+\dots+e_{s}d_{s}=m-t$.
If we choose $\bm{e}=(1,\dots,1)\in{\mathbb{N}}^{s}$, we obtain the classical definition of a $(t,m,s)$-net as given in Definition 1. For $(t,m,s)$-nets, we have the propagation rule that a $(t,m,s)$-net in base $b$ is also a $(v,m,s)$-net in base $b$ for any integer $v$ with $t\leq v\leq m$. However, with the above definition of $(t,m,\bm{e},s)$-nets, we do not have this propagation rule. In [8], the authors provided a revised definition of $(t,m,\bm{e},s)$-nets which ensures the above mentioned propagation rule. In this section, however, we work with the original definition provided by Tezuka in [13].
We note that all $(t,m,s)$-nets are also $(t,m,\bm{e},s)$-nets, however for certain values of $\bm{e}$, we can obtain a lower $t$ value for the corresponding $(t,m,\bm{e},s)$-net. In particular, for column-reduced digital nets, we can find certain examples where at least for some choices of $\bm{e}$, the reduced net retains the original quality parameter $t$. Let us give some examples.
Example 1.
Let $b=2,s=2,m=4$ and consider the $(0,4,2)$-net derived from the Sobol’ sequence, given by the generating matrices
$C_{1}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&1\end{pmatrix},\quad C_{2}=\begin{pmatrix}1&1&1&1\\ 0&1&0&1\\ 0&0&1&1\\ 0&0&0&1\end{pmatrix}.$
Let $w_{1}=0$ and $w_{2}=1$, then the resulting column reduced digital net is a $(1,4,2)$-net according to Theorem 1. However, for $\bm{e}=(e_{1},e_{2})$ chosen such that $(e_{1}d_{1},e_{2}d_{2})=(4,0)$ is the only solution to the equation $e_{1}d_{1}+e_{2}d_{2}=4$, we obtain a column reduced net that is still a $(0,4,\bm{e},2)$-net. This is because the only elementary intervals $J$ that satisfy the conditions in Definition 4 are of the form
$J=[a_{1}/2^{4},(a_{1}+1)/2^{4})\times[0,1).$
Thus, the net property depends only on the first coordinates in the point set ${\mathcal{P}}$, and since we do not set any columns of $C_{1}$ to zero, i.e., $w_{1}=0$, the resulting column reduced net is a $(0,4,\bm{e},2)$-net. Some concrete choices for $\bm{e}$ with the above property are $(e_{1},e_{2})=(2,3)$, or $(e_{1},e_{2})=(1,e_{2})$ where $e_{2}>4$, or $(e_{1},e_{2})=(4,e_{2})$ where $e_{2}=3$ or $e_{2}>4$.
In general, given a $(t,m,s)$-net ${\mathcal{P}}$ derived from a $(t,s)$-sequence and reduction indices $0=w_{1}\leq w_{2}\leq\dots\leq w_{s}$, for $\bm{e}=(e_{1},\dots,e_{s})=(m-t,k,\dots,k)$, where either $1<k<m-t$ with $\gcd(k,m-t)=1$ or $k>m-t>0$, the only solution to the equation $\sum_{j=1}^{s}e_{j}d_{j}=m-t$ is $\bm{d}=(1,0,\dots,0)$. Thus, the column reduced net $\widetilde{{\mathcal{P}}}$ is an $(m-t,m,\bm{e},s)$-net for the above choice of $\bm{e}$.
One might also consider digital $(t,m,s)$-nets generated by $C_{1},\dots,C_{s}$ where the $C_{j}$ for $2\leq j\leq s$ are derived from a $(t,s)$-sequence but $C_{1}$ is not necessarily derived from a digital sequence, since we usually choose the first reduction index $w_{1}=0$. In this case, one could perhaps find more choices of $\bm{e}=(e_{1},\dots,e_{s})$ such that the corresponding column reduced digital net is a $(t,m,\bm{e},s)$-net, depending on the reduction indices $w_{2},\dots,w_{s}$.
A general and complete analysis of reduced $(t,m,\bm{e},s)$-nets could be an interesting subject for future research.
@@ 4.1 A reduced matrix product algorithm
Remark 5.
Let us consider mappings $\bm{\phi}:[0,1]^{s}\rightarrow{\mathbb{R}}$ of the form $\bm{\phi}({\bm{x}})=(\phi_{1}(x_{1}),\ldots,\phi_{s}(x_{s}))$ that we apply simultaneously to all sample points of a given digital net. In this case, we can easily adapt Algorithm 1 such that we obtain a reduced net with points transformed by $\bm{\phi}$, but do not change the order of the computation time outlined in Remark 4. Such an adaption is useful when considering the case $D={\mathbb{R}}^{s}$, as pointed out in the introduction.
@@ 4.2 Error analysis
In this section, we restrict ourselves to the case $D=[0,1]^{s}$, such that we do not need to transform the sample points ${\bm{x}}_{k}$ before applying the corresponding QMC rule. In many applications of quasi-Monte Carlo, one considers so-called weighted function spaces such as weighted Sobolev or weighted Korobov spaces (see, e.g., [3, 5, 6]). The idea of studying weighted function spaces goes back to the seminal paper [12] of Sloan and Woźniakowski. The motivation for weighted spaces is that in many applications, different coordinates or different groups of coordinates may have different influence on a multivariate problem. To give a simple example, consider numerical integration of a function $f:[0,1]^{s}\to{\mathbb{R}}$, where